OpenStudy (anonymous):

solve for x 7^x7^4=7^7

OpenStudy (lgbasallote):

is that (7^x) (7^4) = 7^7? then it's (7^x+4) = 7^7 x+4 = 7 x = 7 - 4 x = 3